# Hits to Higgs: Reconstruction-Free Higgs Classification from Raw LHC Detector Data Using Higgsformers

### v2

@@ 6 Conclusions
Our results demonstrate that the Higgsformer model, trained solely on inner tracker hits, can effectively separate $t\bar{t}H$ from $t\bar{t}$ events, achieving a test ROC AUC of 0.792 at zero pileup. Performance is likely limited by the size of the training dataset and the degree of model optimization.
@@
While object-level models (based on the full detector) still outperform Higgsformer (which uses only inner tracker inputs) in this study, our results show that modern attention-based architectures can extract meaningful event-discriminative information directly from raw hit-level data. This represents a promising step toward reconstruction-free classification in high energy physics.

### v3

@@ 6 Conclusions
Our results demonstrate that the Higgsformer model, trained solely on inner tracker hits, can effectively separate $t\bar{t}H$ from $t\bar{t}$ events, achieving a test ROC AUC of 0.855 at zero pileup for our best model. Performance is likely limited by the size of the training dataset and the degree of model optimization.
@@
While object-level models still outperform Higgsformer for b-tagging efficiencies $f>0.4$, our central conclusion is that non-trivial event-level discrimination can be learned directly from raw inner tracker hits under detector-consistent conditions. These results demonstrate for the first time that modern attention-based architectures can extract meaningful discriminative information directly from hit-level data. This represents a promising step toward hit-level classification in high energy physics, i.e. approaches that do not require explicit intermediate reconstructed objects as inputs.